OpenStudy (superhelp101):

Find a value of k such that 1/2 is a root of 2x^2+11x =-k how do i do this?

OpenStudy (anonymous):

Can you find the roots of quadratic equation by Quadratic Formula?

OpenStudy (anonymous):

Firstly add k to both the sides..

OpenStudy (superhelp101):

yes i can

OpenStudy (superhelp101):

can you find the roots of quadratic equation by Quadratic Formula? i think i can do that

OpenStudy (superhelp101):

2x^2+11x+k

OpenStudy (anonymous):

Just go with that..

OpenStudy (superhelp101):

go with what lol?

OpenStudy (anonymous):

Go with that means go with the quadratic formula, use it and find the roots, do upto that..

OpenStudy (superhelp101):

umm i am kinda confused. but thank you here maybe @iambatman can help?

OpenStudy (anonymous):

Who will write "=0" with that?? Note it..

OpenStudy (superhelp101):

ok

OpenStudy (anonymous):

Trying to avoid calculations so that Batman can do that for you?? Clever..!!

OpenStudy (superhelp101):

why??

OpenStudy (superhelp101):

huh?

OpenStudy (superhelp101):

ohh no!

OpenStudy (superhelp101):

of course not! i want to know how to do the problem! that is why i said "how do i do this?" on my question

OpenStudy (anonymous):

But in comments you are avoiding to do that.. :P

OpenStudy (superhelp101):

thats because i am really confused! :P

OpenStudy (anonymous):

I said to just start it by quadratic formula, just do it, if you are wrong, we will sort out that..

OpenStudy (anonymous):

If you are confused, then you must be confused somewhere, at some place, tell me that place buddy.. :P

OpenStudy (superhelp101):

but, but is this the quadratic formula that you mean??? 2x^2+11x+k

OpenStudy (superhelp101):

that is were i am confused :/ quadratic formula

OpenStudy (anonymous):

Leave it.. I tell from the beginning..

OpenStudy (superhelp101):

ok!

OpenStudy (anonymous):

Firstly you will add k both the sides, because, the general quadratic equation is : \(ax^2 + bx + c = 0\), this "=0" form in right side you have to achieve.. You can get that by adding k both the sides.. So, you will come to : \(2x^2 + 11 + k = 0\) Right?

OpenStudy (superhelp101):

yep!!!

OpenStudy (anonymous):

Now, using quadratic formula, : \[x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

can you tell what is a, b and c here by comparing with the given equation??

OpenStudy (superhelp101):

yeah i can i see what you are doing you are factoring

OpenStudy (anonymous):

It is a kind of factoring, but in actual I am finding out values of x..

OpenStudy (superhelp101):

oh right

OpenStudy (anonymous):

So, here you will get values of x as: \[x = \frac{-11 \pm \sqrt{121 - 8k}}{4}\]

OpenStudy (anonymous):

Here, a = 2, b = 11 and c = k

OpenStudy (superhelp101):

yes i understand that! :)

OpenStudy (anonymous):

Look you have got + and - sign in between, so for one root, you can take anyone.. Suppose I am taking -ve sign: So I have: \(\large x = \frac{-11 - \sqrt{121 - 8k}}{4}\)

OpenStudy (anonymous):

And you are given with that one root is 1/2, so just equate them: \[\frac{-11 - \sqrt{121 - 8k}}{4} = \frac{1}{2}\]

OpenStudy (anonymous):

Now, you need to just solve it for \(k\)..

OpenStudy (superhelp101):

but how do i do that?

OpenStudy (anonymous):

Multiply by 4 firstly both the sides.. Then, Add 11 both the sides.. Do this and show me where you reach..

OpenStudy (superhelp101):

omg can't i just plug in 1/2 into the equation to get k???

OpenStudy (anonymous):

Yes, you can very well do that... :)

OpenStudy (superhelp101):

yayay! but anyway thank you so much for taking you time for me to understand this concept!! :D

OpenStudy (anonymous):

So, what you got for k by doing that?

OpenStudy (superhelp101):

-6